# Copper delocalization leads to ultralow thermal conductivity in chalcohalide CuBiSeCl2

### v1

@@ D. Thermal transport properties
Heat capacity ($C_{V}$) is $3Nk_{B}$ in a typical solid at high temperature (the Dulong–Petit limit), where $N$ is the number of particles and $k_{B}$ is the Boltzmann constant. Here in Fig. 3(a), with temperature going higher than 300 K, the $C_{V}$ given out by DFT calculation is very close to the theoretical value $3Nk_{B}$ = 0.295 J g-1 K-1. Also, this value of CuBiSeCl2 is even lower than 0.36 J g-1 K-1 of liquid-like thermoelectric Cu2Se 26 and a little higher than 0.156 J g-1 K-1 of renowned PbTe 27.
@@
The extent of anharmonicity is typically measured by the Grüneisen parameter, illustrated in Fig. 3(c). Below 1.5 THz, there is an intertwinement between the acoustic and optical phonon branches at 300 K. This is closely linked to the Cu and Bi atoms and the bonding hierarchy. Moreover, the total Grüneisen value is 1.50, close to PbTe value of 1.45 31, indicating CuBiSeCl2 exhibits substantial anharmonicity.
@@
We evaluate the 3ph scattering channels involving three acoustic phonons (AAA, purple), two acoustic phonons combining with one optical phonon (AAO, red), and one acoustic phonon combining with two optical phonons (AOO, orange), as shown in Fig. 3(e). The scattering process of the AAO mode is found to be as strong as the AOO mode below 1 THz, meaning that the AAO and AOO phonons play dominant roles in heat transport. Also, as temperature increases to 700 K in Fig. 3(e) the scattering process of AAA mode becomes much stronger compared to other modes, the maximum scattering rates are changed from around $10^{-4}$ to $10^{-3}$ ps-1, which significantly reduce the $\kappa_{L}$ of CuBiSeCl2 at higher temperatures.
Figure 3: Microscopy heat transport parameters of CuBiSeCl2. (a) Specific heat capacity at constant volume ($C_{V}$). (b) Squared phonon group velocities ($v^{2}$) in the harmonic approximation and anharmonic renormalization approximation at 300 K. (c) Grüneisen parameter. (d) Weighted phonon scattering phase space of 3ph and 4ph at 300 K and 700 K. (e) Phonon scattering rates of 3ph at 300 K and 700 K. (f) Phonon lifetime of at 300 K. The red horizontal line represents the Wigner limit, expressed as $\tau_{\text{threshold}}=[\Delta\omega_{\text{avg}}]^{-1}$, where $\Delta\omega_{\text{avg}}$ is the average phonon inter-band spacing 32. $\Delta\omega_{\text{avg}}$ is determined by $\omega_{\text{max}}/3N_{\text{at}}$, with $N_{\text{at}}$ being the number of atoms per unit cell and $\omega_{\text{max}}$ being the highest frequency. Phonons with lifetimes longer than $\tau_{\text{threshold}}$ primarily affect thermal conductivity via population mechanisms, while shorter lifetimes contribute to wavelike tunneling. The black curve shows the Ioffe-Regel limit, $\tau_{\text{IR}}=2\pi/\omega$, in which $\omega$ is the phonon frequency. Phonons lifetime exceeding $\tau_{\text{IR}}$ can be described by the Wigner transport equation 33, 15, 22.
Recently, a seminal concept is introduced known as the Wigner limit in time, which has been applied to distinguish phonons into various thermal transport regimes, primarily consisting of particle-like propagation and wavelike tunneling 32, 25, 34. The Wigner limit in time is characterized as $\tau_{\text{threshold}}=3N_{\text{at}}/\omega_{\text{max}}$, where $N_{\text{at}}$ represents the number of atoms in the unit cell, and $\omega_{\text{max}}$ denotes the maximum phonon frequency. According to this criterion, phonons with lifetimes exceeding $\tau_{\text{threshold}}$ exhibit particle-like behavior and are the main contributors to the $\kappa_{p}$ from phonon population. In contrast, those with shorter lifetime exhibit wavelike tunneling behavior, primarily contributing to the $\kappa_{c}$ from coherence phonons.
@@
Moreover, quartic anharmonicity not only induces substantial phonon frequency shifts but also results in pronounced 4ph scattering. We obtained the average $\kappa_{p,SCPH}^{3,4ph}$ values of 4.24, 0.58, and 0.32 W m-1 K-1 at 50, 300, and 700 K, respectively, shown in Fig. 4(b). These values represent reductions of 17%, 33%, and 42% compared to $\kappa_{p,SCPH}^{3ph}$, highlighting the impact of 4ph scattering 13, 18, 35. This result confirms that phonon hardening and 4ph scattering, driven by quartic anharmonicity, have counteractive effects on $\kappa_{L}$.
@@
Figure 4: (a) A schematic illustration demonstrating the prediction accuracy of $\kappa_{L}$ across different hierarchical thermal transport methods, including the HA+3ph, SCPH+3ph, SCPH+3,4ph, and SCPH+3,4ph+$\kappa_{c}$, where 3ph indicates only three-phonon scattering is included and 3,4ph means both three-phonon and four-phonon scattering are considered. HA means harmonic approximation, and SCPH is short for self-consistent phonon. (b) $\kappa_{p}$ of CuBiSeCl2 using HA+3ph, SCPH+3ph, SCPH+3,4ph approaches. (c) $\kappa_{p}$ of CuBiSeCl2 compare with $\kappa_{c}$. (d) $\kappa_{L}$ of CuBiSeCl2 using BTE method and MD method compare with experimental values 3.
@@
Nevertheless, the coherence contributions are not negligible as shown in Fig. 4(c). By further considering the off-diagonal terms of heat flux operators, the total $\kappa_{L}$ significantly enhanced by incorporating both $\kappa_{p,SCPH}^{3,4ph}$ and $\kappa_{c}$ 25. The ratio $\kappa_{c}/\kappa_{p}$ (red star) is greater than 0.5 at 500 K and exhibits a pronounced increase with temperature. Specifically, in the $a$-axis direction, the ratio of $\kappa_{c}/\kappa_{p}$ (blue cubic) reaches 0.94 at 700 K, which is 1.7 times larger than the 0.55 observed at 300 K. This observation suggests that the contribution of $\kappa_{c}$ is equally important as $\kappa_{p}$ in crystalline CuBiSeCl2 especially in $a$-axis and high temperature. The dominant role of coherence contribution in thermal transport was also observed in other complex compounds such as Cs2AgBiBr6 33, Cu12Sb4S13 36, CsPbBr3 37, TlAgI2 15, La2Zr2O7 32, Tl9SbTe6 22, and even simple cubic CsCl 35.
@@
We further use the HNEMD method to assess the intrinsic $\kappa_{L}$ of CuBiSeCl2 19, 21, 38. Since the MD method includes full-order lattice anharmonicity larger than those lattice anharmonicity considered in BTE up to fourth-order, the results are relatively smaller than those obtained from BTE. Fig. 4(d) reveals that the average $\kappa_{p,SCPH}^{3,4ph}$ at 300 K is 0.584 W m-1 K-1, larger than 0.466 W m-1 K-1 from HNEMD. The latter HNEMD result is a little larger than the experimental $\kappa_{L}$, which is 0.275 W m-1 K-1 at 300 K. This could result from their non-dense sample, with a density of only 88% 3, 39. The result suggests that the higher than fourth-order lattice anharmonicity could be crucial for capturing the ultralow $\kappa_{L}$ of complex compounds such as CuBiSeCl2.
@@ Conclusion
In summary, this study explores the thermal transport properties of CuBiSeCl2 by utilizing a combination of self-consistent phonon theory based on machine learning potentials, first-principles calculations, and a unified approach of the Wigner transport equation to account for both coherent and population contributions to thermal transport. The neuroevolution potentials achieve a level of accuracy comparable to DFT in predicting energies and atomic forces. The temperature-dependent force constants reveal a significant phonon hardening effect over nearly the entire frequency spectrum, indicating strong lattice anharmonicity in CuBiSeCl2. This leads to ultralow $\kappa_{L}$ of 0.81 W m-1 K-1 at 300 K. By using the HNEMD method, where full-order lattice anharmonicity is considered, the $\kappa_{L}$ decreases further to 0.47 W m-1 K-1, closely aligning with experimental measurements 3, 39.
$\kappa_{c}/\kappa_{p}$ exceeds 0.5 above 400 K, indicating that the off-diagonal components of heat flux operators play a significant role in the thermal transport of CuBiSeCl2. Most strikingly, we trace ultralow $\kappa_{L}$ of CuBiSeCl2 to the delocalization of loosely bound copper atoms. Our work may shed light on further suppressing $\kappa_{L}$ and seek for promising thermoelectrics by adjusting bonding strength, localization, and delocalization of electrons 26, 40.
@@
The authors gratefully acknowledge discussions with Jiongzhi Zheng. We acknowledge the support from the National Natural Science Foundation of China (No.12104356 and No.52250191), and the Fundamental Research Funds for the Central Universities. This work is sponsored by the Key Research and Development Program of the Ministry of Science and Technology (No.2023YFB4604100). We also acknowledge the support by HPC Platform, Xi’an Jiaotong University.

### v3

@@ C. Thermal transport properties
Heat capacity ($C_{V}$) is $3Nk_{B}$ in a typical solid at high temperature (the Dulong–Petit limit), where $N$ is the number of particles and $k_{B}$ is the Boltzmann constant. Here in Fig. 3(a), with temperature going higher than 300 K, the $C_{V}$ given out by DFT calculation is very close to the theoretical value $3Nk_{B}$ = 0.295 J g-1 K-1. Also, this value of CuBiSeCl2 is even lower than 0.36 J g-1 K-1 of liquid-like thermoelectric Cu2Se 41 and a little higher than 0.156 J g-1 K-1 of renowned PbTe 42.
@@
The extent of anharmonicity is typically measured by the Grüneisen parameter, illustrated in Fig. 3(c). Below 1.5 THz, there is an intertwinement between the acoustic and optical phonon branches at 300 K. This is closely linked to the Cu and Bi atoms and the bonding hierarchy. Moreover, the total Grüneisen value is 1.50, close to PbTe value of 1.45 46, indicating CuBiSeCl2 exhibits substantial anharmonicity.
@@
We evaluate the 3ph scattering channels involving three acoustic phonons (AAA, purple), two acoustic phonons combining with one optical phonon (AAO, red), and one acoustic phonon combining with two optical phonons (AOO, orange), as shown in Fig. 3(e). The scattering process of the AAO mode is found to be as strong as the AOO mode below 1 THz, meaning that the AAO and AOO phonons play important roles in reduce heat transport. Also, as temperature increases to 700 K in Fig. 3(e) the scattering process of AAA mode becomes much stronger compared to other modes, the maximum scattering rates are changed from around $10^{-4}$ to $10^{-3}$ ps-1, which significantly reduce the $\kappa_{L}$ of CuBiSeCl2 at higher temperatures.
Figure 3: Microscopy heat transport parameters of CuBiSeCl2. (a) Specific heat capacity at constant volume ($C_{V}$). (b) Squared phonon group velocities ($v^{2}$) in the harmonic approximation and anharmonic renormalization approximation at 300 K. (c) Grüneisen parameter. (d) Weighted phonon scattering phase space of 3ph and 4ph at 300 K and 700 K. (e) Phonon scattering rates of 3ph at 300 K and 700 K. (f) Phonon lifetime of at 300 K. The red horizontal line represents the Wigner limit, expressed as $\tau_{\text{threshold}}=[\Delta\omega_{\text{avg}}]^{-1}$, where $\Delta\omega_{\text{avg}}$ is the average phonon inter-band spacing 47. $\Delta\omega_{\text{avg}}$ is determined by $\omega_{\text{max}}/3N_{\text{at}}$, with $N_{\text{at}}$ being the number of atoms per unit cell and $\omega_{\text{max}}$ being the highest frequency. Phonons with lifetimes longer than $\tau_{\text{threshold}}$ primarily affect thermal conductivity via population mechanisms (green region), while shorter lifetimes contribute to wavelike tunneling (blue region). The black curve shows the Ioffe-Regel limit, $\tau_{\text{IR}}=2\pi/\omega$, in which $\omega$ is the phonon frequency. Phonons lifetime exceeding $\tau_{\text{IR}}$ can be described by the Wigner transport equation 48, 33, 39.
Recently, a seminal concept is introduced known as the Wigner limit in time, which has been applied to distinguish phonons into various thermal transport regimes, primarily consisting of particle-like propagation and wavelike tunneling 47, 26, 49. The Wigner limit in time is characterized as $\tau_{\text{threshold}}=3N_{\text{at}}/\omega_{\text{max}}$, where $N_{\text{at}}$ represents the number of atoms in the unit cell, and $\omega_{\text{max}}$ denotes the maximum phonon frequency. According to this criterion, phonons with lifetimes exceeding $\tau_{\text{threshold}}$ exhibit particle-like behavior and are the main contributors to the $\kappa_{p}$ from phonon population. In contrast, those with shorter lifetime exhibit wavelike tunneling behavior, primarily contributing to the $\kappa_{c}$ from coherence phonons.
@@
Moreover, quartic anharmonicity not only induces substantial phonon frequency shifts but also results in pronounced 4ph scattering. We obtained the average $\kappa_{p,SCPH}^{3,4ph}$ values of 4.24, 0.58, and 0.32 W m-1 K-1 at 50, 300, and 700 K, respectively, shown in Fig. 4(b). These values represent reductions of 17%, 33%, and 42% compared to $\kappa_{p,SCPH}^{3ph}$, highlighting the impact of 4ph scattering 31, 36, 50. This result confirms that phonon hardening and 4ph scattering, driven by quartic anharmonicity, have counteractive effects on $\kappa_{L}$.
@@
Figure 4: (a) A schematic illustration demonstrating the prediction accuracy of $\kappa_{L}$ across different hierarchical thermal transport methods, including the HA+3ph, SCPH+3ph, SCPH+3,4ph, and SCPH+3,4ph+$\kappa_{c}$, where 3ph indicates only three-phonon scattering is included and 3,4ph means both three-phonon and four-phonon scattering are considered. HA means harmonic approximation, and SCPH is short for self-consistent phonon. (b) $\kappa_{p}$ of CuBiSeCl2 using HA+3ph, SCPH+3ph, SCPH+3,4ph approaches. (c) $\kappa_{p}$ of CuBiSeCl2 compare with $\kappa_{c}$. (d) $\kappa_{L}$ of CuBiSeCl2 using BTE method and MD method compare with experimental values 4.
@@
Nevertheless, the coherence contributions are not negligible as shown in Fig. 4(c). By further considering the off-diagonal terms of heat flux operators, the total $\kappa_{L}$ significantly enhanced by incorporating both $\kappa_{p,SCPH}^{3,4ph}$ and $\kappa_{c}$ 26. The ratio $\kappa_{c}/\kappa_{p}$ (red star) is greater than 0.5 at 500 K and exhibits a pronounced increase with temperature. Specifically, in the $a$-axis direction, the ratio of $\kappa_{c}/\kappa_{p}$ (blue cubic) reaches 0.94 at 700 K, which is 1.7 times larger than the 0.55 observed at 300 K. This observation suggests that the contribution of $\kappa_{c}$ is equally important as $\kappa_{p}$ in crystalline CuBiSeCl2 especially in $a$-axis and high temperature. The dominant role of coherence contribution in thermal transport was also observed in other complex compounds such as Cs2AgBiBr6 48, Cu12Sb4S13 51, CsPbBr3 52, TlAgI2 33, La2Zr2O7 47, Tl9SbTe6 39, and even simple cubic CsCl 50.
@@
We further use the HNEMD method to assess the intrinsic $\kappa_{L}$ of CuBiSeCl2 27, 29, 16. Since the MD method includes full-order lattice anharmonicity larger than those lattice anharmonicity considered in BTE up to fourth-order, the results are relatively smaller than those obtained from BTE. Fig. 4(d) reveals that the average $\kappa_{p,SCPH}^{3,4ph}$ at 300 K is 0.584 W m-1 K-1, larger than 0.466 W m-1 K-1 from HNEMD. The latter HNEMD result is a little larger than the experimental $\kappa_{L}$, which is 0.275 W m-1 K-1 at 300 K. This could result from their non-dense sample, with a density of only 88% 4, 53, 54, 37. The result suggests that the higher than fourth-order lattice anharmonicity could be crucial for capturing the ultralow $\kappa_{L}$ of complex compounds such as CuBiSeCl2.
@@ Conclusion
In summary, this study explores the thermal transport properties of CuBiSeCl2 by utilizing a combination of self-consistent phonon theory based on machine learning potentials 55, 16, 17, first-principles calculations, and a unified approach of the Wigner transport equation to account for both coherent and population contributions to thermal transport. The neuroevolution potentials achieve a level of accuracy comparable to DFT in predicting energies and atomic forces. The temperature-dependent force constants reveal a significant phonon hardening effect over nearly the entire frequency spectrum, indicating strong lattice anharmonicity in CuBiSeCl2. This leads to ultralow $\kappa_{L}$ of 0.81 W m-1 K-1 at 300 K. By using the HNEMD method, where full-order lattice anharmonicity is considered, the $\kappa_{L}$ decreases further to 0.47 W m-1 K-1, closely aligning with experimental measurements 4, 53.
$\kappa_{c}/\kappa_{p}$ exceeds 0.5 above 400 K, indicating that the off-diagonal components of heat flux operators play a significant role in the thermal transport of CuBiSeCl2. Most strikingly, we trace ultralow $\kappa_{L}$ of CuBiSeCl2 to the delocalization of loosely bound copper atoms. Our work may shed light on further suppressing $\kappa_{L}$ and seek for promising thermoelectrics by adjusting bonding strength, localization, and delocalization of electrons 41, 56, 57..
@@
The authors gratefully acknowledge discussions with Jiongzhi Zheng. We acknowledge the support from the National Natural Science Foundation of China (No.52250191), and the Fundamental Research Funds for the Central Universities. This work is sponsored by the Key Research and Development Program of the Ministry of Science and Technology (No.2023YFB4604100). We also acknowledge the support by HPC Platform, Xi’an Jiaotong University.
Data Availability
The structure data of CuBiSeCl2 that we used can be found in Supplemental Material 37. Other data is available from the authors upon reasonable request.